### v2

# active distribution network DER distributed energy resource

Abstract
This paper presents a grid-aware probabilistic approach to compute the aggregated flexibility at the grid connection point (GCP) of active distribution networks to allow the participation of DERs in ancillary services (AS) markets. Specifically, an optimal power flow (OPF) method is used to compute the aggregated capability for the provision of multiple AS at the GCP of an ADN over the full AS provision horizon. The proposed method extends current approaches for the time-coupled flexibility aggregation, allowing to consider multiple services simultaneously, ensure cost-effectiveness of the used DERs and handle uncertainties in a probabilistic way. The disaggregation of the total flexibility into individual DERs power flexibilities accounts for the operational costs associated to the provision of different services. This ensures cost-effectiveness while maximizing the value of the aggregated flexibility, assuming known service prices. Empirical uncertainty sets are obtained to achieve a predefined coverage of the probability distribution in line with recent developments in the Nordic AS markets. Finally, a feeder-decomposition approach is proposed to ensure applicability to realistic distribution networks with a large number of buses. Case studies show the effectiveness of the method, highlight the importance of network constraints and time-coupling and illustrate its applicability to realistic distribution systems.
Index Terms:
Flexibility Aggregation, Ancillary Services, Optimal Power Flow, Robust Optimization
I Introduction
The large penetration of renewable energy generation and the increasing electrification of various processes change the paradigm of electricity system operation. Stochastic fluctuations of electricity generation, combined with concentrated peak demand, lead to more uncertainty in the operation of the electricity system. This is reflected in the increasing amounts of balancing services and the associated prices for secondary frequency regulation (aFRR) [1]. In response to these evolutions, the authors of [2] stated that ”It is essential to take advantage of the opportunity to harness the valuable and increasing amount of resources at the distribution level for providing services for the overall benefit of the power system.”. New regulations allowing DERs to participate in energy and flexibility markets provide wholesale market benefits and new revenue streams for DERs owners and operators [3]. At the same time, DERs, aggregated in virtual power plants, provide a viable alternative to conventional generation for cost-effective provision of grid balancing services [4]. As a consequence, DERs are competing with conventional service providers, such as fuel-based generators, which are to be phased out in accordance with international greenhouse gas targets (e.g. [5]). Additionally, most DERs are connected at the distribution level. This leads to increased loading of distribution systems and impacts the overall steady state and dynamic performance of the bulk power system [6]. Stronger cooperation between distribution system operators and transmission system operators and appropriate strategies to make DER flexibility available for the provision of grid services at the transmission level are required.
I-A DSO-TSO Interaction
This section briefly highlights the importance of DSO-TSO coordination, before Section I-B reviews existing approaches for the coordination mechanism of interest, flexibility aggregation. Readers interested specifically in coordination mechanisms are referred to [3]-[7] and the references therein. The power systems community has largely acknowledged the need for closer cooperation between TSOs and DSOs [6]. As the volume of services provided by DERs increases, DSOs must be actively involved to avoid issues at the distribution level [7]. Furthermore, by only optimizing the operation at the distribution level, solutions that are locally optimal may not lead to a global optimum of the whole system. A broad overview of the main approaches for DSO-TSO coordination is provided in [7]. The authors differentiate between three models based on the responsibilities of both system operators. In a first option, the DER bids are directly transmitted to the TSO, who selects and activates the required bids accounting for the DSO constraints. The second approach considers that the DSO first validates the bids based on its own operational constraints, before the TSO selects the preferred bids. Finally, in the ’DSO-managed’ model, the DSO validates and aggregates the DER bids, the TSO selects the required aggregated flexibility and the DSO sends the resulting activation commands. Assessing the aggregated flexibility in ADNs is beneficial for both the planning and operation [3].
I-B Flexibility aggregation approaches
We consider a DSO-TSO cooperation where the DSO aggregates the DER capabilities to offer flexibility to the TSO. The relevant problem is to quantify and explicitly determine the available flexibility, while accounting for network and resource constraints and stochastic boundary conditions.  [^1] In this respect, the considered problem is different than approaches presented in [8] and [9], where the operation of a distribution system under uncertainty is considered and works such as [10], where clearing of energy and flexibility markets is considered while accounting for DERs. In this section, existing approaches for flexibility aggregation are discussed and the remaining issues are identified. A first distinction between the methods reported in literature can be made based on the type of algorithm used [11]. The first group consists of approaches based on Monte Carlo (MC) simulations. Here, a significant number of scenarios are considered, each corresponding to a feasible actuation of the flexible resources. For each scenario, the corresponding power exchange at the GCP is obtained. If all constraints are satisfied, the corresponding power exchange is labeled as a feasible point within the PQ plane. By performing a large number of simulations, an accurate mapping of the power capability can be made. An example of such an approach is presented in [12]. However, as expected, the authors report that many scenarios are needed to guarantee an accurate representation of the flexibility set, leading to large computation times, and that additionally it is hard to capture the extreme points of the capability map. Recent developments (e.g. [13]) consider this problem and propose a Bayesian learning approach to alleviate the computational challenges. However, they still rely on an initial outer-box approximation of the flexibility set obtained through optimization.
On the other hand, optimization-based approaches can be divided in a group of methods that attempt to compute the exact power aggregation set at the GCP and a group of methods that compute an approximate set, which can typically be described more easily. We further differentiate existing methods based on four important features of the flexibility aggregation sets: grid-awareness, time-coupling, cost-awareness and wether they account for uncertainty. As an example, [14] proposes an approach to compute the exact aggregated power flexibility set of multiple DERs without accounting for grid constraints. They also show that the number of constraints may be too large to solve the exact problem and have instead proposed a set of approximate models. Recognizing this issue, [15] proposes to cluster DERs based on the similarity of their capabilities and obtain an inner approximation through a specific type of polytope. The DERs in each cluster can then be exactly aggregated. Alternatively, in [16] a ”geometric prototype” is constructed by selecting a subset of the equations forming the exact power aggregation set, keeping geometrical features of the exact model while reducing the computational complexity. Approaches attempting to identify the exact aggregated flexibility set, while simultaneously satisfying the grid constraints, also exist. In this case, a set of OPF problems is typically solved. For instance, in [17] a set of non-approximated non-convex OPF problems are solved to obtain points on the aggregated capability curve in the PQ plane until a convergence criterion based on the distance between the obtained points is satisfied. This approach also allows to include cost constraints. Similar approaches are also proposed in [18], [19], [20] and [21]. Explicit computation of flexibility costs for these approaches is proposed in [22], but costs are computed a posteriori for identified points of the flexibility map and consider a single time step. These approaches only allow to compute the aggregation set for a single time step as the solved OPF problems cannot account for time coupling. To alleviate this problem, [23] proposes a network-informed data driven approach based on a classifier labeling samples to obtain an approximation allowing for temporal coupling.
While the above methods can obtain flexibility areas closer to the exact set, they are only applicable when neglecting grid constraints or considering a single time step or have to resort to approximations to make the problem tractable. Therefore, methods seeking convex inner approximations of the exact flexibility set are desirable. An approach for the estimation of a robust aggregation set is presented in [24], where a linear power flow model is used. However, only single time steps are considered and DERs costs are not taken into account. In [25] a grid-aware approach, using a fixed-point linearization, is presented to compute the aggregate power flexibility over multiple time steps by approximating it as a hyperbox. However, uncertainty of stochastic prosumption is not accounted for. The authors attempt to model the costs, but only the one of the base trajectory, around which flexibility is offered, is included. In [26] the authors show that the flexibility aggregation set obtained through a hyperbox approximation is very conservative and that a larger one can be obtained with an inner ellipsoidal approximation. Using the same fixed-point linearization, the authors present tractable reformulations to obtain the maximum-volume ellipsoid under affine and quadratic disaggregation policies, while accounting for uncertainty through ellipsoidal uncertainty sets at each time step. An approximate approach decomposing the flexibility aggregation set in a ”virtual battery” and ”virtual generator” is presented in [27]. Uncertainty is accounted for through chance constraints assuming Gaussian probability distributions suitably transformed to deterministic constraints. The authors claim better results compared to the ellipsoidal inner approximation. The same authors extend this in [28] with a piecewise fitting of the aggregated cost function for the different time steps. However, the cost function is computed by minimizing the operational cost for different points on the time-decoupled aggregation sets and fitting an affine function. Therefore, it does not account for the time-coupling of the flexibility and its influence on the costs. An alternative approach to the approximate aggregation problem is proposed in [29]. By solving the disaggregation problem for samples selected in the flexibility aggregation space and labeling feasible points a convex ellipsoidal classifier is found. This ellipsoid is used as a surrogate solution space within which a polytope mapping the aggregated flexibility is determined. Uncertainty is accounted for when constructing the surrogate set. All the works above consider the flexibility aggregation problem to be the mapping of a single flexibility commodity at the ADN GCP. However, even though all flexibility services require power exchange, the provision of multiple ones (e.g. primary frequency regulation (FCR) and aFRR) have different implication on time-coupling and, therefore, the feasibility of an aggregation set. This is demonstrated in works regarding the optimal provision of AS using battery energy storage systems, such as [30]. Furthermore, in the works considering uncertainty, forecast errors are either modeled robustly, leading to overly conservative solutions, or through chance constraints related to individual time steps. This does not accurately represent the impact of stochastic variables. Finally, although some works consider costs, none of these methodologies enforce cost-effectiveness of flexibility in an appropriately time-coupled manner. To summarize, none of the above approaches consider simultaneously network constraints, time-coupling and cost of the flexibility and the impact of uncertainty. The provision of multiple AS is also not considered in state-of-the-art aggregation approaches. Therefore this paper makes the following contributions:
  1. 1.
    We propose a multi-service flexibility aggregation approach based on (quadrant)-ellipsoidal sets.
  2. 2.
    We integrate in the aggregation problem explicit constraints modeling the disaggregation and ensuring any selected flexibility is cost-effective.
  3. 3.
    Uncertainty is accounted for in a probabilistic way through joint ellipsoidal uncertainty, following the P90 requirement [31] for the availability of power flexibility.
II Problem Statement
The problem considered in this work is the flexibility aggregation at the GCP of DERs within ADNs with generic topologies (i.e. either meshed or radial) to provide multiple AS at the transmission level. To accurately model the available flexibility, it is essential to consider both the capabilities of DERs and the constraints of the network. This ensures the aggregated flexibility does not adversely impact the operation of the distribution system. Additionally, the impact of uncertain prosumption and the time-coupling of the DERs flexibility capabilities is highly relevant. Consider the general formulation (1), where $p_{0}$ is the vector representing the power exchanged at the GCP over all time steps, $\Omega_{p_{0}}$ is the set of feasible $p_{0}$ values, $p$ and $q$ are stacked vectors containing the vectors of controllable active and reactive power injections [^2] Note that in this formulation, the controllable injections are assumed to be known, with a capability curve that is constrained by the uncertainties. over all time steps and $\zeta$ is the stacked vector collecting the uncertainty drivers $\zeta_{t}\in\mathcal{U}_{t}$ over all time steps.
$\max_{\Omega_{p_{0}},p,q}$ $\quad J(\Omega_{p_{0}}):\forall\zeta_{t}\in\mathcal{U}_{t},t=1..T$ (1a)
s.t. $G(p,q)+b(\zeta)=p_{0},\quad L(p,q)\leq z(\zeta)$ (1b)
$J$ represents the flexibility maximizing objective, $G$ and $b$ represent the mapping between respectively the controllable resources and the uncertainty drivers and the slack power, while $L$ and $z$ represent the network and resource constraints. The goal of (1) is to determine the flexibility-maximizing set $\Omega_{p_{0}}$ of power exchanges at the GCP. Any point within this set should be feasible for any realization of the uncertainty drivers $\zeta_{t}$. Directly obtaining such a set is in general untractable, even when neglecting grid constraints and using linear DER models [32]. The authors of [26] propose a convex, tractable approach to determine the maximum power flexibility of an ADN by restricting the set of power exchange trajectories at the GCP to an ellipsoid, following (2) with $e$ the center and $E$ describing the axes and shape. Any point within this ellipsoid represents a vector of power exchanges at the ADN for the considered horizon. The use of (2) makes $\Omega_{p_{0}}$ explicit.
[…]
In Section II-A we summarize important results from [26] before we extend them in the next sections to determine the flexibility potential for the provision of multiple AS. To this end, Section II-C shows the aggregation of multiple AS, Section II-E introduces cost-effectiveness, Section II-D proposes a probabilistic approach to integrate uncertainties and Section II-F extends the method for multiple feeders.
II-A Network model and power aggregation
The presented approach is applicable to any linear power flow model (e.g. as presented in [33]). In this work, the model introduced by [34] is used. Network constraints are linearized using a fixed-point equation, which can be interpreted as a linear interpolation of local power flow linearizations around two operating points, leading to a good global approximation. The linear grid model [^3] An explicit expression for the current magnitudes is not provided in [34] as they can be written as second order cone constraints using the real and complex parts of the line currents. However, a linear model for the branch current magnitudes is obtained similarly as for the nodal voltage magnitudes. is presented in (3) with $p$ and $q$ representing the active and reactive controllable power injections.
$p_{0}=G\begin{bmatrix}p\\ q\end{bmatrix}+b,\;|v|=[K_{p}K_{q}]\begin{bmatrix}p\\ q\end{bmatrix}+w,\;|i|=[J_{p}J_{q}]\begin{bmatrix}p\\ q\end{bmatrix}+d$ (3)
The coefficient matrix $G$ represents a linear mapping for the power balance, with constant coefficient $b$. Finally, $K_{p}$, $K_{q}$ and $J_{p}$, $J_{q}$ represent the linear coefficients mapping the power injections to the nodal voltage magnitudes $|v|$ and the branch current magnitudes $|i|$ respectively, with $w$ and $d$ the corresponding constant terms. Using the grid model (3), the network constraints are written as (4).
$v_{min}\leq|v|\leq v_{max},0\leq|i|\leq I_{max}$ (4)
By using linear approximations for all resource constraints, together with a linear power flow model and restricting the flexibility set to an ellipsoid, the flexibility-maximizing problem can be rewritten as (5) [26] .
$\max_{E,e,p}$ $\quad\textrm{log}(\textrm{det}(E)):\forall||\xi||\leq 1,\forall||\zeta_{t}||\leq 1,t=1..T$ (5a)
$\textrm{s.t.}\quad G$ $\begin{bmatrix}p\\ q\end{bmatrix}+B\zeta+b_{0}=E\xi+e,\;W\begin{bmatrix}p\\ q\end{bmatrix}\leq Z\zeta+z_{0}$ (5b)
$B$ and $b_{0}$ and $Z$ and $z_{0}$ are the affine mappings from $\zeta$ to $b$ and $\zeta$ to $z$ respectively. $W$ models the inequalities, including (4). Equality constraints should be eliminated when possible in robust optimization problems [35]. To this end, the authors of [26] exploit the structure of the problem. By writing $p=B_{1}x+B_{2}y$, with the columns of $B_{1}$ being an orthogonal basis for $G$ and the columns of $B_{2}$ spanning the null-space of $G$, equality constraints are eliminated. Finally, to make this problem tractable, an affine policy (6c) is introduced, with linear dependencies $K$ and $L_{t}$ on the position on the ellipsoid and the uncertainties respectively and a constant $\gamma$. With these reformulations and the introduction of $W_{1}=WB_{1}$, $W_{2}=WB_{2}$ $D=GB_{1}$, the problem is given by (6).
$\max_{E,e,K,\gamma,L_{t}}\quad\textrm{log}(\textrm{det}(E))$ (6a)
s.t. $W_{1}D^{-1}(E\xi+e-(B\zeta+b_{0}))+W_{2}y$
$\leq$ $Z\zeta+z_{0}\quad\forall||\xi||\leq 1,\forall||\zeta_{t}||\leq 1,t=1..T$ (6b)
$y=K\xi+\gamma+\sum_{t}L_{t}\zeta_{t}$ (6c)
To reformulate the robust constraints in (6) the worst case for each over the ellipsoidal aggregation and uncertainty sets is considered for each constraint individually. This results in the deterministic problem (7) where $\Theta_{t}$ represents the affine dependence on $\zeta_{t}$, $\nu$ collects the constant offsets, $w_{1,i}$, $w_{2,i}$ are the ith row of $W_{1}$, $W_{2}$ and $\alpha_{i}$ is an auxiliary variable.
$\max_{E,e,K,\gamma,L_{t}}\textrm{log}(\textrm{det}(E))$ (7a)
s.t. $\forall i=1..m\quad\alpha_{i}+w_{1,i}D^{-1}e+w_{2,i}\gamma-\nu_{i}\leq 0$ (7b)
[…]
II-B Resource Constraints
The power flexibility aggregated at the GCP is provided by the controllable DERs in the ADN. Although their character is typically truly non-linear [36], resource capabilities may be accurately described through linear constraints, as shown for example in [37]. This allows a linear representation of all the constraints limiting the flexibility provision, as required to eliminate the equality constraints in (5). Here we only note the constraints for a BESS, but this can be generalized to other DERs. The particularity of BESSs is that they introduce time-coupling, meaning the flexibility aggregation problem can not be solved for each $t$ separately. The battery active and reactive power $p_{b},q_{b}$ and state of energy (SOE) constraints $\forall t$ are:
$p_{b}^{min}\leq$ $p_{b}\leq p_{b}^{max},\;SOE^{min}\leq SOE(t)\leq SOE^{max}$ (8a)
$q_{b}^{min}\leq$ $q_{b}\leq q_{b}^{max},\;SOE(t+1)=SOE(t)-p_{b}(t)\$(8b)
II-C Multiple services provision
One of the main contributions of this work is to determine the flexibility capacity for multiple AS simultaneously. Given that all services are provided by the same DERs, meaning the different service compete for the same power flexibility, the objective must account for the value of the different services. To quantify this value, the different service prices are used. This reflects the willingness of system operators to pay and thus the service value for the system. An important feature of this formulation is that it allows to consider the aggregation of flexibility for different services in a non-hierarchical fashion. The multi-service problem (9) is written as an extension of (5). Note that the GCP power for the different services $p_{0}^{s}$ is formulated here for active power services, as these are the services where time-coupling (e.g. due to limited energy storage) is most relevant. The formulation is however applicable for both active and reactive power services.
$\max_{p^{s},\Omega_{p^{s}_{0}}}$ $\quad\sum_{s}J^{s}(\Omega_{p^{s}_{0}})$ (9a)
s.t. $\forall\zeta\in\mathcal{U},\forall s\in\mathcal{S}\quad\sum_{s}(W^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix})\leq Z\zeta+z_{0}$ (9b)
$G^{s}\begin{bmatrix}p^{s}\\ q^{s}\end{bmatrix}+B^{s}\zeta+b^{s}_{0}=p^{s}_{0}\quad p^{s}_{0}\in\Omega_{p^{s}_{0}}$ (9c)
This objective does not consider the operational costs of DERs. Section II-E details how the DER costs are accounted for and why they are not included in the objective function. In the remainder of this work, we focus on active power services provided at the GCP. We now specify the different types of services that are considered, where an important distinction must be made between symmetric (e.g. FCR) and asymmetric services (e.g. aFRR). All services must be modeled separately as they might have different prices at the same time step, they have different power and energy requirements and their provision might cause different costs. Symmetric services require that for each flexibility bid with a power rating $p_{0s}$, any realization between $p_{0s}$ and $p_{0s}$ must be feasible. For asymmetric services, a power setpoint between 0 and $p^{{}^{\prime}}_{0s}$ or $p^{{}^{\prime\prime}}_{0s}$ and 0 can be requested. For each service, the determined flexibility capacity at the GCP represents the range of power set points that can be realized by the ADN. A robust representation of the exchanged power is therefore required. Additionally, corresponding to standard market practices, the different services for which flexibility is offered, must be available independently. For linear grid and resource models this problem can be written in a linear way by using the superposition principle to sum the power contributions for the different services independently. A separate equality constraint can then be written for each service, linking the service slack power to the disaggregated contributions of the controllable resources. Geometrically, stacking different services can be seen as determining the maximum volume ellipsoids, weighted with the service prices, for which the Minkowski sum lies within the polytope representing the problem’s constraints. Each service is related to the different constraints and the constraint satisfaction depends on all the services independently, reflecting independent selection and activation of the bids. Symmetric services, ranging from $p^{t}_{0s}$ to $p^{t}_{0s}$ at each time step, can be naturally represented as ellipsoids, with some restrictions on the orientation. For asymmetric services, bids can be either for positive or negative power exchange. Any point between zero and the selected bid must be feasible. A natural representation for these services, based on the ellipsoidal capability aggregation, comes in the form of quadrant ellipsoids, where a positive quadrant ellipsoid will be used to represent a positive asymmetric service. Below the used aggregation sets for the different services are explained.
  • For symmetric services, the flexibility is represented as an ellipsoid. Additionally, the ellipsoid must be restricted to be symmetric around all the axes. Without this restriction, the ellipsoids do not lead to valid FCR ranges. Figure 1 illustrates this condition. The red ellipsoid is not symmetric around the axes, meaning that the capacity for upward regulation does not necessarily match the capacity for downward regulation, as illustrated by the red dots. By restricting the ellipsoids to be symmetric around all axes, we obtain ellipsoids of the blue type, for which the blue dots illustrate the matching up/down regulation. With this restriction, $E^{sm}$ becomes diagonal and $e^{sm}$ vanishes. Finally, without loss of generality, the elements of $E^{sm}$ are restricted to be positive as they represent the maximum symmetric power capacity that can be offered at any time step and both positive and negative power exchanges are included through the realization of $\xi$. Symmetric services are then described by (10).
[…]
    $\begin{bmatrix}p^{sm}\\ q^{sm}\end{bmatrix}=(B_{1}^{sm}D^{{sm}^{-1}}E^{sm}+B_{2}^{sm}K^{sm})\xi$ (10b)
    (a) Symmetric Ellipsoid
    (b) Asymmetric Ellipsoid
    Figure 1: Ellipsoidal flexibility aggregation: 2D example.
  • For asymmetric services, with upwards regulation, the flexibility is represented as a positive quadrant ellipsoid, restricted by (11). The restrictions imposed for the symmetric services are also enforced for the asymmetric services. This is needed to represent the services as quadrant ellipsoids, as off-diagonal terms may lead to negative power injections for a $\xi$ in the positive quadrant.
[…]
    $\begin{bmatrix}p^{ap}\\ q^{ap}\end{bmatrix}=(B_{1}^{ap}D^{{ap}^{-1}}E^{ap}+B_{2}^{ap}K^{ap})\xi$ (11b)
  • For asymmetric services with downwards regulation, the flexibility set becomes a negative quadrant ellipsoid (12).
[…]
    $\begin{bmatrix}p^{an}\\ q^{an}\end{bmatrix}=(B_{1}^{an}D^{{an}^{-1}}E^{an}+B_{2}^{an}K^{an})\xi$ (12b)
Note that for all the services, both the slack power $p^{0,s}$ and the controllable injections $p^{s},q^{s}]^{T}$ are fully linear in $\xi$ with no constant component, as the service activation is uncertain and the flexibility set must include the zero activation option.
Next to the power flow caused by the offered flexibility, power flows caused by the baseload are included, namely, the power exchanged at the GCP that is not linked to any AS. In this work, we consider three use cases for the base power, referring to the different TSO-DSO coordination approaches. The base power can be uncontrollable, in which case it simply reflects the power flowing at the GCP as a consequence of the uncontrollable prosumption. Alternatively, the base power can be controlled, for example by a centralized DSO who optimizes the expected energy cost of operating the ADN. In the following, this case is referred to as the ’baseload control’. Both these cases are occurrences of the third DSO-TSO coordination scheme, where the DSO validates and transmits the aggregated bids to the TSO. Finally, we also consider the case of ’self-dispatching’. In this case, the base power at different time steps is restricted to an ellipsoid and controllable resources are activated to balance the uncertain prosumption and limit the uncertainty at the GCP. This can be seen as a special case of the fourth coordination scheme, where the DSO is responsible for its own imbalances and only the remaining flexibility can be offered to the TSO. The modeling of baseload power for these cases is detailed here below.
  • In the self-dispatching case, the goal is to minimize the uncertainty on the baseload perceived at the GCP. The baseload power is represented as an ellipsoid using the approach introduced Section II-A. This naturally reduces the uncertainty on the base power exchange as any point within the ellipsoid must be feasible for any realization of the uncertainty, meaning a larger ellipsoid would reserve a larger portion of the DERs flexibility. This leads to the following set of possible baseload powers, as in (6).
[…]
    $p^{b}=(B_{1}^{b}D^{{b}^{-1}}(E^{b}\xi+(e^{b}-b(\zeta)))+B_{2}^{b}(K^{b}\xi+\gamma^{b}+\sum_{u}L^{u}\zeta^{u})$
  • For the controllable baseload power, two new variables $p_{0}^{b}$ and $y^{b}$ are introduced. These represent the expected slack power and controllable injections (14). In this case only the expected value $p^{b}_{0}$ is optimized.
    $p^{b}=B_{1}^{b}D^{b^{-1}}(p_{0}^{b}-b(\zeta))+B_{2}^{b}y^{b}$ (14)
  • In the uncontrolled baseload case, the base slack power is the net result of the prosumption at the different nodes. This can be seen as a special case of the baseload control case, where $p^{b}=0$ and $y^{b}=0$
II-D Probabilistic Constraint Satisfaction
The available power flexibility in ADNs depends on the stochastic prosumption, as it influences both the network constraints and the capabilities of DERs. Failing to account for these uncertainties leads to an overestimation of the ADN’s power flexibility. Alternatively, accounting for all possible realizations of these uncertain power injections and robustly ensuring the flexibility capacity is available can be overly conservative. Certain grid operators have already recognized this will be an issue when more flexibility is supplied by DERs [31]. To allow DERs with uncertain capabilities to participate in the AS markets, the grid operator allows actors to offer flexibility services as long as the forecast available capacity is available at least 90% of the time [38] (i.e. according to the P90 requirement). To this end, the stochastic prosumption is handled in a probabilistic way in this work. Note that this availability requirement should be interpreted as a joint chance constraint as merely enforcing the flexibility capacity to be present with 90% probability at each time step will lead to more frequent overbidding. As joint chance constraints are generally hard to solve, we determine a robust uncertainty set for the uncertainty drivers that guarantees a certain coverage $1-\epsilon$ of the empirical probability distribution. For uncertainty drivers with gaussian distributions, exact ellipsoidal uncertainty sets can be obtained [39]. Additionally, we consider the computation of ellipsoidal uncertainty sets with a $1-\epsilon$ coverage for the empirical probability distribution of the available data. A comparison of these approaches with robust ellipsoidal sets covering all scenarios is presented in Section III-A. In this work, we assume a single uncertainty driver $\zeta^{u}$ suffices to model the stochastic fluctuations of all individual injections of a certain type. With the ellipsoidal uncertainty sets, the stochastic injections can be formulated as $p^{u}=\bar{p}^{u}+\Delta\zeta^{u}\>\forall u\in\mathcal{U}$, with $\Delta$ a linear mapping from the uncertainty drivers to the stochastic power injections [^4] Consider for example PV plants. The power can be approximated as $P^{cap}*GHI$, where the global horizontal irradiance ($GHI$) uncertain.. Note that ellipsoidal uncertainty sets can be obtained for each time step separately, considering as stochastic variables $\zeta_{t}^{u}$ with $\zeta_{t}^{u}||\leq 1$ at time $t$ as was done in [26]. However, to guarantee a coverage of the joint probability distribution and jointly satisfy the problem constraints, a single ellipsoidal uncertainty set for $\zeta=[\zeta^{u1},\zeta^{u1}...\>\zeta^{u|\mathcal{U}|}]$ with $\zeta||\leq 1$ should be considered for all uncertainty drivers over the full horizon.
II-E Ensuring Cost-Effectiveness
To realistically represent the flexibility ADNs can provide, it is crucial to take into account the operational costs for the provision of the different services. Accounting for these ensures the flexibility is cost-effective and thus ”economically feasible” [40]. Therefore, we derive a cost-effectiveness condition that can be formulated in a convex way under some mild assumptions on the cost functions of the DERs. The costs for providing the different services can be obtained through the affine disaggregation policy used to map the flexibility provided at the GCP to the individual DERs. The controllable injections can be obtained from (10b), (11b) and (12b).
$p^{s}=(B_{1}^{s}D^{s^{-1}}E^{s}+B_{2}^{s}K^{s})\xi$ (15)
A sufficient condition for the convexity of the problem is that the costs for the DERs must be linear with respect to the allocated power capacity. This is a reasonable assumption as we show in Appendix -A for the example of BESSs. For linear cost functions, the flexibility cost becomes:
$C^{s}=c^{s^{T}}|(B_{1}^{s}D^{s{-1}}E^{s}+B_{2}^{s}K^{s})\xi|$ (16)
where $c^{s}$ is a cost vector containing the linear cost coefficients for the DERs for the provision of service $s$. This expression can be further simplified, as the flexibility capabilities are represented as (quadrant) ellipsoids, which means we can force all controllable injections to have the same sign. This means that they all contribute to the flexibility.
$B_{1}^{s}D^{s}{{}^{-1}}E^{s}+B_{2}^{s}K^{s}\geq 0$ (17)
The costs and benefits can then be equivalently reformulated as in (18) for service ellipsoids with positive elements.
$C^{s}=c^{s^{T}}(B_{1}^{s}D^{s{-1}}E^{s}+B_{2}^{s}K^{s})|\xi|,\quad B^{s}=g_{s}^{T}E^{s}|\xi|$ (18)
Constraining the benefits to be higher than the costs ensures cost-effectiveness (19) for any set of activated services.
$g^{s^{T}}E^{s}-c^{s^{T}}(B_{1}^{s}D^{s{-1}}E^{s}+B_{2}^{s}K^{s})\geq 0$ (19)
One might desire to directly optimize the net benefit (19). However, two key issues appear. First, the net benefit depends on the selected flexibility meaning the real benefit is expressed as:
$B^{s}_{net}=\sum_{s}(g^{s^{T}}E^{s}-c^{s^{T}}(B_{1}^{s}D^{s{-1}}E^{s}+B_{2}^{s}K^{s})|\xi|^{s}$ (20)
Given that any $\xi||\leq 1$ can be selected, the realized benefit is not known apriori. Secondly, optimizing for the best case as an example would lead to the maximization of a norm, making the problem non-convex. Therefore the objective in (21) is selected as a proxy to maximize the flexibility value. Here,the quantity of flexibility offered at each time step is represented by the diagonal elements of $E^{s}$. Denoting the service prices at time steps $t$ as $p^{s}_{t}$, the objective becomes:
$J^{s}=\sum_{t}p^{s}(t)E^{s}(t,t),\quad J=\sum_{s}J^{s}$ (21)
Finally, to avoid biasing the energy available in the baseload case, two constraints are added.
$\sum_{t}(p^{b}_{0}-b)=0$ (22)
$\sum_{t}(c^{0}(t)p^{b}_{0}(t)-\sum_{r}|c^{r}(t)^{T}|p^{b}(t)|)\geq 0$ (23)
The first one (22) enforces the integral of the difference of baseload power with respect to the expected net load to be zero. The second (23) ensures the power balancing accounts for operational costs, where $r$ indexes the different DERs. Including these cost-effectiveness considerations, the final problem is given by Problem (24). The impact of this controlled baseload on the problem constraints is modeled through (24m) and the influence of the uncontrollable prosumption and corresponding adjustment of the base power is included in (24g). Note that the subscript $u$ refers to the number of uncertainty sets to ensure the formulation remains generic. In the previous section, based on the work from [26], uncertainty sets were defined per timestep. In Section II-D we show the benefits of a different approach. The influence of all the services and the selected baseload control is combined in (24h) and (24i) where $e^{0}$ and $\gamma$ disappear for the baseload case. We introduce auxiliary variables $\alpha^{i}$ for each constraint and $\alpha^{i}_{s}$ to constrain the impact of the different services on each constraint. Reformulating the problem with robust constraints and restricting the aggregated flexibility set to (quadrant) ellipsoids for all services yields the multiservice aggregation problem (24).
$\max_{E^{s},K^{s}}\quad\sum_{s}\sum_{t}p^{s}(t)E^{s}(t,t)$ (24a)
[…]
$w^{ap}_{1,i}D^{ap^{-1}}E^{ap}+w^{ap}_{2,i}K^{ap}\leq\epsilon_{i}^{ap}$ (24c)
[…]
$w_{1,i^{an}}D^{an^{-1}}E^{an}+w^{an}_{2,i}K^{an}\geq\epsilon_{i}^{an}$ (24e)
[…]
$\sum_{u}w^{0}_{2,i}L^{0}_{u}-\theta_{u,i}\leq\lambda_{i}$ (24g)
$\alpha^{i}\geq\sum_{s}\alpha_{s,i}+\sum_{t}\lambda_{t,i}$ (24h)
$\alpha_{i}+\beta_{i}+w^{0}_{1,i}D^{-1}e^{0}+w^{0}_{2,i}\gamma^{0}-\nu_{i}\leq 0$ (24i)
$g^{s^{T}}E^{s}-c^{s^{T}}(B_{1}^{s}D^{s{-1}}E^{s}+B_{2}^{s}K^{s})\geq 0$ (24j)
$B_{1}^{s}D^{s}{{}^{-1}}E^{s}+B_{2}^{s}K^{s}\geq 0$ (24k)
$\textrm{if self-dispatching}:$
[…]
if baseload:
$w^{0}_{1,i}D^{0^{-1}}p_{0}+w^{0}_{2,i}y^{0}\leq\alpha^{0}_{i}\quad$ (24m)
$\sum_{t}(p^{b}_{0}-b)=0$ (24n)
$\sum_{t}(c^{0}(t)p^{b}_{0}(t)-\sum_{r}|c^{r}(t)^{T}|p^{b}(t)|)\geq 0$ (24o)
II-F Multifeeder Aggregation
For scalability, an additional step allowing to aggregate multiple feeders connected to the same primary substation is developed. In this case, the aggregated flexibility is first computed for each feeder separately, by solving (24). In a second step, using the obtained aggregation sets, a further aggregation set for all the considered feeders is obtained by solving a problem similar to the single feeder aggregation. The contributions of the different feeders replace the contributions of the DERs and the resource capabilities are replaced by the feeder ellipsoids $E_{f}^{s}$ for each service $s$, leading to second order cone (SOC) constraints. As all DERs constraints are embedded in the feeder ellipsoids, only the transformer rating constraints need to be added in this multifeeder aggregation problem. Given that any set of flexibility bids must be feasible irrespective of the total baseload power, the transformer rating is adjusted based on the range of possible baseload powers computed in the feeder aggregation problem. For consistency, the same objective is used as in the single feeder problem, requiring the positivity constraint (17) to model cost-effectiveness. This leads to the following formulation, where (25b) reflects the feeder flexibilities and (25c) ensures the transformer, with upper and lower power limits $p^{max}_{transfo}$ and $p^{min}_{transfo}$, is not overloaded.
$\max_{E^{s},K^{s}}\quad\sum_{s}\sum_{t}p^{s}_{t}E^{s}(t,t)$ (25a)
[…]
$b=[p^{max}_{transfo}-p^{max}_{base},p^{min}_{transfo}+p^{min}_{base}]^{T}$ (25d)
Since all feeders are fed by the same primary substation transformer, the individual aggregation problems can be solved separately by enforcing a robust slack voltage constraint at each feeder. Indeed, by allowing only solutions that satisfy all feeder constraints for voltages in between $1-\delta$ and $1+\delta$ pu, the problems can be decoupled while maintaining overall feasibility. This is done by duplicating the voltage and current network constraints for the extreme values of the considered slack voltage range. The total flexibility to be provided by the multifeeder ADN is then disaggregated first between the different feeders, according to the multifeeder disaggregation policy and then further disaggregated within each feeder, according to their individual policies.
III Results
This section demonstrates the approach for a number of test cases. First results for a the IEEE33 benchmark system are shown. Then the approach is extended to multiple feeders, interconnected to a common primary substation, showing its effectiveness when aggregating the flexibility of realistic distribution systems. Finally the impact of network constraints is discussed. Common inputs for the service prices and the uncertainty sets are used. Input data was obtained from the Réseau de Transports d’Electricité [41] (for the prices), Deutscher Wetterdienst [42] (for the PV production) and historical residential data from [43] for the loads. Figure 2 shows the inputs used for the case studies.
(a) Load (top) and PV scenarios (bottom).
(b) Prices for service provision and operation costs.
Figure 2: Prosumption and costs data, from [41], [42] and [43].
Figure 3: Uncertainty sets from a) a fitted multivariate gaussian distributions, b) robust with respect to all scenarios, c) 1-$\epsilon$ coverage rate with respect to all scenarios.
III-A Comparison of the Ellipsoidal Uncertainty Sets
In Figure 3, we show the boundaries of the ellipsoidal sets obtained in different ways. In all cases, the considered ellipsoidal uncertainty sets offer a coverage of 90%. First, shown in yellow, we show the robust ellipsoids, considering all the possible uncertainty realizations based on the prediction set. Obtaining robust ellipsoidal sets is a well-known problem [44] and consists of determining the minimum volume ellipsoid containing all points in given set. Next, we consider ellipsoids obtained in a probabilistic way, guaranteeing a $90\%$ coverage with respect to the empirical distributions. A first approach to obtain such a set, shown in blue, is based on a multivariate Gaussian distribution, obtained by fitting the prediction scenarios. As the load and PV uncertainties are not truly normally distributed, this leads to an overly conservative uncertainty set. A second approach assumes the set of past observations is representative for the future realizations. This allows to obtain probabilistic uncertainty sets using a heuristic. By computing the mutual distance between all points and selecting the one with the lowest maximum distance to the $1-\epsilon$ points with the smallest distance to this point. There is no guarantee this will yield the ellipsoid with the smallest volume for a certain coverage $1-\epsilon$, but it leads to significant improvements compared to the robust case. This is shown in green in the figure. Representing the ellipsoids is not possible due to the large number of dimensions. Instead we show the range of possible values at all time steps for the different approach, (i.e. the extreme points in all dimensions). The range of prediction scenarios and historical realizations used to obtain the uncertainty sets are also shown.
III-B Benchmark Distribution System
The IEEE33 benchmark feeder containing $1.4MWp$ of PV injections and $2.7MWp$ load is selected as a benchmark system. Four controllable DERs are considered, in this case all BESSs with a total energy capacity of $24MWh$ and a total power capacity of $6MW$. The BESS capacities were obtained by increasing the storage capacity until the self-dispatching version of the problem became feasible, allowing a comparison between the different aggregation versions. Time-coupled uncertainty sets are used for both PV and load. Specifically, the uncertainty sets obtained by considering the empirical $90\%$ coverage were used. Figure 4(a) shows the results when only the base power trajectory is controlled in a stacked plot. In this case more flexibility can be offered at the transmission level as uncertainties only impact potential congestions and other grid constraints but do not reserve energy or power capacity of the flexible DERs at the distribution level. aFRR is selected due to the higher service prices in this case, with no FCR provision for both the baseload and self-dispatching due to the lower prices.
(a) Aggregated flexibility offered at the feeder GCP.
(b) Contribution of the resources to the offered flexibility.
Figure 4: Aggregated flexibility for the IEEE33 (baseload).
Figure 5 shows the results for the self-dispatching case, in which the flexibility is first used to compensate local uncertainties before services are offered to the upper-level system. Only upward aFRR is offered due to the limited flexibility. Note that the GCP baseload power shown in Figure 5(a) does not represent the total local balancing energy budget as both upwards and downwards capacity are reserved. Much less flexibility is available in the self-dispatching case, compared to the baseload case, as a large energy budget is reserved to balance deviations of the stochastic prosumption from the baseload. To show the utilization of the DERs storage assets and validate the aggregation sets, we run MC simulations, where for each selected set of flexibility actuations, we sample the prosumption from the considered uncertainty sets. The BESS SOEs trajectories are shown in Figure 6 for the self-dispatching case, showing the storage limits are satisfied. Additionally, note that the SOE stays far from the limits. This is due to the energy capacity reserved to balance the uncertain prosumption, for which only a limited number of realizations where used in the a posteriori sampling. Other quantities such as the ADN’s nodal voltages and branch currents are also validated as shown later in Section III-D.
(a) Aggregated flexibility offered at the feeder GCP.
(b) Contribution of the resources to the offered flexibility.
Figure 5: Aggregated flexibility for the IEEE33 (selfdispatching).
Figure 6: BESSs SOE trajectories from MC simulations.
III-C Flexibility Potential in Realistic Distribution Systems
To demonstrate the multifeeder approach presented in Section II-F, we solve the aggregation problem for five similar ADNs obtained from the synthetic networks presented in [45], with the network IDs given in Table I. The loading and PV integration for all networks are added in Table I together with the controllable DERs characteristics. The considered voltage range at the slack node of the individual ADNs or feeders is set to $\delta=0.02\>pu$. Figure 7 shows the results for two cases of the transformer rating at the full network GCP. The service price of FCR was considered double the value reported in Figure 2(b) to show the provision of all services. Reducing the transformer rating to $27MW$ limits the available flexibility. For the symmetric FCR service, this leads to a reduction in both directions although the transformer constraint is only binding for additional power consumption.
(a) Aggregated flexibility offered at each feeder GCP.
(b) Aggregated ADN flexibility for two transformer ratings (top: $30MW$, bottom: $27MW$). The dotted line represents the tightened transformer rating on both.
Figure 7: Flexibility for a multifeeder ADN (baseload).
Network IDs from [45] ID 1 ID 2 ID 4 ID 7 ID 8
Number of nodes 23 19 27 24 17
Load Rating [kW] 3205 2531 5035 2589 2911
PV Capacity [kW] 1380 3085 4496 9236 4456
BESS power capacity [kW] 1800 4400 5200 6600 3600
BESS energy capacity [kWh] 3600 8800 10400 13200 7200
Number of loads 18 10 17 8 9
Number of PV injections 5 9 10 16 8
Number of BESSs 3 3 4 4 2
TABLE I: Characteristics of ADNs for the multifeeder case.
For the application of the proposed aggregation approach, the computational complexity of the method is also relevant. Due to the complex nature of the problem, the computation time increases unacceptably when directly considering large distribution networks. The multifeeder approach allows to decompose large distribution networks per feeder to keep the problem tractable. For the two cases on the IEEE33 benchmark case, the computations times were 30 and 18 minutes for the self-dispatching and the baseload case respectively. For the multifeeder aggregation, the mean and max time per feeder were 24 and 45 minutes respectively. We recall that thanks to the robust voltage constraints, each feeder problem can be solved in parallel. The problem combining the feeders takes a few seconds. Solving the multifeeder problem in one go (without the disaggregation in feeders) was intractable on the adopted computer (a MacBook M1 Max with 32GB RAM).
III-D Importance of grid-aware flexibility aggregation.
This section shows how the network constraints impact the aggregated flexibility. To this end, the aggregated flexibility for the IEEE33 network is compared with the case where all network constraints are neglected. To create additional congestions highlighting the importance of modeling the network constraints, we set all line ampacity limits to a third of their nominal value. For the baseload case, removing network constraints means uncertainty plays no role when considering only BESSs as flexible DERs as non of the constraints are affected by uncertainty drivers. Therefore the full flexibility can then be used for TSO requirements. Figure 8(a) shows the aggregated flexibility obtained with and without grid constraints. When the grid constraints are not accounted for, more flexibility can be offered at time steps with higher forecast service prices (cfr Figure 2(b)). Due to the time-coupling, more flexibility is offered in the grid-aware approach for some time steps. Note that the additional flexibility offered in the grid-unaware case is the result of an overestimation of the available flexibility and is not deliverable due to the violation of grid constraints. Also here, no FCR is provided. Figure 8(b) shows histograms of the line loading for the lines connecting the BESSs for the grid-aware and grid-unaware cases. Currents are computed a posteriori for a set of MC simulations where both the flexibility exchanged and the uncertain prosumption are varied. The frequent violations of the line constraints in the grid-unaware case show the importance of accounting for grid constraints when determining the available flexibility. The results also show that the exact currents slightly differ from the ones predicted by the linear model in the optimization. Therefore in some cases, the true currents may exceed the ampacity limits, as shown for example in Figure 8(b) for BESS2, but in general the limits hold.
(a) Flexibility with and without network constraints.
(b) Currents for the grid-aware and grid-unaware cases.
Figure 8: Impact of network constraints.
IV Conclusions
In this work, we develop a new method to compute the aggregated flexibility of DERs hosted in ADNs for the provision of multiple AS. The proposed method accounts for the uncertainty of prosumption in a probabilistic way through ellipsoidal uncertainty sets and ensures cost-effectiveness by accounting for DERs operation costs. The value of flexibility from a system operators point of view is maximized, assuming the service prices are known. Finally, we demonstrate the applicability of the method to realistic distribution grids by adding robust voltage constraints and combining the aggregated flexibility of multiple feeders while accounting for the ADN transformer rating. The proposed framework improves the coordination between DSOs and TSOs by representing the available flexibility for different ancillary services separately. The presented approach represent the capability of DERs at the distribution level to provide flexibility at the transmission level and thus increase the number of market participants. The affine disaggregation policies for each AS further allow to directly control the DERs in accordance with the selected flexibility. Additionally, by integrating the costs of the different resources, the aggregated flexibility can be represented using SOC constraints, which could be directly integrated in conic markets, such as advocated by [46]. Further work will include investigating approaches to integrate uncertainties without the need for pre-computed uncertainty sets.